# Recall, Robustness, and Lexicographic Evaluation

### v5

@@ 3.1. Metric Orientation
We are interested in more precisely defining precision and recall as constructs to be measured in ranking evaluation. Although most evaluation metrics colloquially capture some aspects of both precision and recall, understanding the sensitivity to each remains vague. We can address this vagueness by approaching precision and recall as two extremes of recall requirements. At one extreme, precision as a construct reflects the satisfaction of a user who only needs exactly one relevant item, the minimum amount of retrievable content. We might find this in domains like web search. At the other extreme, recall as a construct reflects the satisfaction of a user who needs every relevant item, the maximum amount of retrievable content. Zobel et al. (2009) refers to this as the totality interpretation of recall, found in many technology-assisted review domains.
Indeed, this perspective is supported by evaluation programs like the TREC Total Recall Track (Roegiest et al., 2015) and patent search (Lupu et al., 2017); and by metrics like ‘position of the last relevant ’ (Zou and Kanoulas, 2020).
@@
Figure 4. Metric orientation of ranking metrics of $10^{5}$ items with $m\in[1\mathrel{{.}\,{.}}\nobreak 15]$ relevant items. The vertical axis reflects, for $m\in[1,15]$, the change in metric value when (3(a)) moving top-ranked item from position 1 to position $n-m+1$ or (3(b)) moving bottom-ranked item from position $m$ to position $n$. The values for TSE (Equation 3) are scaled by the lower and upper bound given a fixed $m$ and therefore apply to any exposure model. See Figure 3 for details. This figure best rendered in color.
@@
Figure 4. Metric orientation of ranking metrics of $10^{5}$ items with $m\in[1\mathrel{{.}\,{.}}\nobreak 15]$ relevant items. The vertical axis reflects, for $m\in[1,15]$, the change in metric value when (3(a)) moving top-ranked item from position 1 to position $n-m+1$ or (3(b)) moving bottom-ranked item from position $m$ to position $n$. The values for TSE (Equation 3) are scaled by the lower and upper bound given a fixed $m$ and therefore apply to any exposure model. See Figure 3 for details. This figure best rendered in color.
In terms of precision orientation, the ordering of metrics follows conventional wisdom. $\mathrm{RR}$ is often used for known-item or other high-precision tasks where the user is satisfied by the first relevant item. Across all values of $m$, $\mathrm{RR}$ dominates other metrics. $\mathrm{NDCG}$, often used for web search evaluation, is next more precision-oriented metric. AP, ‘recall at 1000’ ($\text{R}_{1000}$), and R-precision (RP) all have the same precision-orientation and are dominated by $\mathrm{NDCG}$. Rank-biased precision ($\mathrm{RBP}$) dominates $\mathrm{NDCG}$ and AP once we reach a modest number of relevant items. Both $\mathrm{RR}$ and $\mathrm{RBP}$ are not sensitive to the number of relevant items because neither is a function of $m$.
@@ 3.2. Total Search Efficiency
Although metrics such as AP are often referred to as ‘recall-oriented’, in this section, we focus on metrics that explicitly define recall as a construct.
Such recall metrics for ranking systems come in two flavors. [^5] We exclude set-based metrics used in set-based retrieval or some technology-assisted review evaluation (Cormack and Grossman, 2016; Lewis et al., 2021), since they require systems to provide a cutoff in addition to a ranking. The first flavor of recall metrics measures the fraction of relevant items found after a user terminates their scan of the ranked list. Metrics like $\text{R}_{1000}$ and RP use a model of search depth to simulate how deep a user will scan. We can define exposure and normalization functions for $\text{R}_{k}$ and RP,
@@
Similarly, Zou and Kanoulas (2020) uses ‘position of the last relevant item’ to evaluate high-recall tasks. By contrast, Rocchio (1964) proposed recall error, a metric based on the average rank of relevant items,
@@ 4. Robustness
In the context of a single ranking, we are interested in measuring its robustness in terms of its effectiveness for different possible users who might have issued the same request. [^6] Robustness in the information retrieval community has traditionally emphasized slightly different notions from our ranking-based perspective. For example, the TREC Robust track emphasized robustness of searcher effectiveness across information needs, focusing evaluation on difficult queries (Voorhees, 2004). Similarly, risk-based robust evaluation seeks to ensure that performance improvements across information needs are robust with respect to a baseline (Wang et al., 2012; Collins-Thompson et al., 2013). Meanwhile, Goren et al. (2018) proposed robustness as the stability of rankings across adversarial document manipulations. In the context of recommender systems, robustness has analogously focused on robustness of utility across users (Xu et al., 2020; Wen et al., 2022) and stability of rankings across adversarial content providers (Mobasher et al., 2007). This is related to work in search engine auditing that empirically studies how effectiveness varies across different searchers issuing the same request (Mehrotra et al., 2017). In our work, we define robustness as the effectiveness of a ranking for the worst-off user who might have issued a request.
@@
Underlying our notion of robustness is a population-based perspective on ranking evaluation. Classic effectiveness measures can be interpreted as expected values over different user populations defined by different browsing behavior (Robertson, 2008; Sakai and Robertson, 2008; Carterette, 2011). For example, Robertson (2008) demonstrated that AP can be interpreted as the expected precision over a population of users with different recall requirements. More generally, Carterette (2011) demonstrated this for a large class of metrics. We can contrast this with online evaluation production environments where systems observe individual user behavior and do not need to resort to statistical models to capture different user behavior. So, just as recent work in the fairness literature disaggregates evaluation metrics to understand how performance varies across groups (Mehrotra et al., 2017; Ekstrand et al., 2018, 2022; Neophytou et al., 2022), we can disaggregate traditional evaluation metrics to understand how performance varies across implicit subpopulations of users such users or providers.
@@ 4.1. User Robustness
Given a ranking $\pi$, we would like to measure the worst-case effectiveness over a population of possible user.
@@ 4.1.1. Possible Information Needs
Similarly, in the context of recommendation, the notion of topical relevance can be interpreted as the potential utility—but not the psychological relevance or desirability—of an item to the user. So, while a particular item may be of interest to a user in general, at any one point in time, it may be undesired for any number of reasons. First, the context of a user can affect whether a particular item is desired. For example, in music recommendation, a user may be interested in bluegrass music in many contexts but, while studying, may prefer ambient music. Second, as with information retrieval, the desirability of a previously-consumed relevant item may degrade due to order effects. For example, in music recommendation, consider a user who listened to a specific relevant song $X$. There are two possible effects. On the one hand, satiation (Leqi et al., 2021) means that the user may not want to listen to $X$ immediately again, making it no longer relevant to that user. Or, obsessive listening (Ward et al., 2014; Conrad et al., 2019) means that the user may want to listen to $X$ over again, making other otherwise relevant songs no longer relevant to that user. So, from the perspective of relevance, $\mathcal{R}$ should be considered the union of psychologically relevant items over all possible states of knowledge a user might have when approaching the system. Indeed, multiple authors describe topical relevance as necessary but not sufficient for psychological relevance (Boyce, 1982; Ruthven, 2014; Cooper, 1971). These papers suggest that rather than seeing a ranking system as acting to directly provide psychologically relevant items, it provides topically relevant items that are candidates to be scanned for psychologically relevant items by the user (Boyce, 1982). In light of this discussion, instead of considering any item in $\mathcal{R}$ as definitely satisfying the user’s information need, we consider it as only having a nonzero probability of satisfying the user’s information need.
@@ 4.1.2. Robustness Across Possible Information Needs
In other words, the worst-off user is the one associated with $p_{m}$, the lowest-ranked relevant item. We present a proof in Appendix B.1. This result implies that recall-orientation captures the utility of a ranking for the worst-off user.
@@ 4.2.1. Possible Provider Preferences
Just as with information needs, each ranking consists of exposure of multiple possible providers. Consider the domains like job applicant ranking systems or dating platforms, where each item in the catalog is associated with an individual person. We assume that each relevant provider $i\in\mathcal{R}$ is interested in its cumulative positive exposure in the ranking (Section 2.3.2), $\eta(\pi,\mathcal{R},\{i\})$. In the more general case, providers can possibly be associated with multiple relevant items in $\mathcal{R}$. This might occur if a creator contributes multiple items to the catalog (e.g. multiple songs, videos, documents); or, a provider may aggregate content from multiple individual creators (e.g. publishers, labels). Even if we have metadata attributing groups of items to specific providers or creators, their preferences for exposure of those items may be unobserved. Provider preferences can themselves complex, covering a broad set of commercial, artistic, and societal values (Hadida, 2015; Eikhof and Haunschild, 2006).
@@ 4.2.2. Robustness Across Possible Provider Preferences
Given this definition, we can show that $\operatorname{TSE}$ is equal to the utility of the worst-case provider (proof in Appendix B.1).
@@ 5. Lexicographic Evaluation
In Section 3, we defined recall-orientation from the perspective of users interested finding in the totality of relevant items and then proposed a new metric, $\operatorname{TSE}$, based on this interpretation. In Section 4, we demonstrated how $\operatorname{TSE}$ measures the worst-case utility for multiple definitions of users and providers, connecting it to notions of robustness and fairness, through Rawls’ difference principle. In this section, we will further develop the fairness perspective by combining recent work in preference-based evaluation with classic work in social choice theory, improving the nuance in worst-case analysis and allowing it to be useful as an evaluation tool. We will begin by discussing the practical limitations of $\operatorname{TSE}$ for evaluation (Section 5.1) before developing a preference-based evaluation method derived from social choice theory that generalizes $\operatorname{TSE}$ and improves its practical use (5.2).
@@ 5.1. Low Sensitivity of Total Search Efficiency
Although measuring worst-case performance and, as a result, Rawlsian fairness, $\operatorname{TSE}$ may not satisfy our desiderata for an evaluation method (Section 2.4). To understand why, consider two rankings $\pi$ and $\pi^{\prime}$ for the same request. When comparing a pair of systems, we are interested if defining a preference relation $\pi\succ\pi^{\prime}$. The worst-case preference $\pi\succ_{\text{WC}}\pi^{\prime}$, is defined as,
@@ 5.2. Lexicographic Recall
where $i=\max\{j\in[1\mathrel{{.}\,{.}}\nobreak m]:x_{i}\neq y_{j}\}$. We show a example of this process in Figure 8. This way of comparing rankings generates a total lexicographic ordering over vectors of the same dimensionality and is often used in the fairness literature to address ties when adopting Rawls’ difference principle. [^7] For further discussion of the connection between Rawls’ difference principle and leximin, see (Sen, 1976, 1977; Kolm, 2002; d’Aspremont and Gevers, 1977; Moulin, 2003).
@@
Although operating over $\mathcal{U}$ grounds our evaluation in possible user information needs, scoring and ranking $m^{+}$ subsets of $\mathcal{R}$ can be computationally intractable. So, just as we demonstrated that we only need to inspect the position of the last relevant item to compute $\succ_{\text{WC}}$, we can demonstrate that we only need to compare the rank positions of the relevant items to compute $\pi\succ_{\text{LR}}\pi^{\prime}$ (proof in Appendix B.2) and, therefore,
@@
where $i=\max\{j\in[1\mathrel{{.}\,{.}}\nobreak m]:p_{i}\neq p^{\prime}_{j}\}$. Moreover, although defined as a user-oriented metric, we can also demonstrate that this results in provider leximin as well (proof in Appendix B.2),
@@ 5.3. Sensitivity of Lexicographic Recall
We can demonstrate the higher sensitivity of lexirecall through simulation and analysis of the total space of permutations. In Figure 7, we demonstrated that, for a set of random paired rankings, the number of ties was high for traditional metrics and grew quickly for $\operatorname{TSE}$ as the cutoff $\tilde{n}$ decreased. Figure 7 also includes lexirecall, which exhibits substantially fewer ties than traditional metrics and $\operatorname{TSE}$. Independent of simulation, we are also interested in the probability of a tie over for randomly sampled pairs of complete rankings $\pi,\pi^{\prime}\in S_{n}$ (i.e. $\tilde{n}=n$). We can derive these probabilities (see Appendix D) as functions of $m$, $n$, and any parameters of the metric (e.g. $k$),
@@
To better understand the relationship between these probabilities, we display probabilities of ties for several retrieval depths in Table 3. Although Figure 7 demonstrated that $\operatorname{TSE}$ exhibited poor sensitivity when $\tilde{n}<n$, it is much more sensitive for complete rankings, in part because random complete rankings are less likely to share $p_{m}$ than imputed rankings. Both traditional recall metrics exhibit many more ties, especially as the retrieval depth grows. Amongst methods, lexirecall and AP demonstrate the few or no ties across corpus sizes. Table 4 presents the same results for varying numbers of relevant items. The results are consist with Table 3, where traditional recall metrics exhibit a large number of ties, which decreases slowly with the number of relevant items. By comparison, $\operatorname{TSE}$, lexirecall, and AP show higher sensitivity with a negligible number of ties.

### v6

@@ 3.1. Metric Orientation
We are interested in more precisely defining precision and recall as constructs to be measured in ranking evaluation. Although most evaluation metrics colloquially capture some aspects of both precision and recall, understanding the sensitivity to each remains vague. We can address this vagueness by approaching precision and recall as two extremes of recall requirements. At one extreme, precision as a construct reflects the satisfaction of a user who only needs exactly one relevant item, the minimum amount of retrievable content. We might find this in domains like web search. At the other extreme, recall as a construct reflects the satisfaction of a user who needs every relevant item, the maximum amount of retrievable content. Herlocker et al. (2004) call this the ‘find all good items’ tasks; Zobel et al. (2009) refers to this as the totality interpretation of recall, found in many technology-assisted review domains. Indeed, this perspective is supported by tasks like recommender systems for scholarly literature reviews; evaluation programs like the TREC Total Recall Track (Roegiest et al., 2015) and patent search (Lupu et al., 2017); and by metrics like ‘position of the last relevant ’ (Zou and Kanoulas, 2020).
@@
Figure 5. Metric orientation of ranking metrics of $10^{5}$ items with $m\in[1\mathrel{{.}\,{.}}\nobreak 15]$ relevant items. The vertical axis reflects, for $m\in[1,15]$, the change in metric value when (5(a)) moving top-ranked item from position 1 to position $n-m+1$ or (5(b)) moving bottom-ranked item from position $m$ to position $n$. The values for TSE (Equation 3) are scaled by the lower and upper bound given a fixed $m$ and therefore apply to any exposure model. See Figure 4 for details. This figure best rendered in color.
In terms of precision orientation, the ordering of metrics follows conventional wisdom. $\mathrm{RR}$ is often used for known-item or other high-precision tasks where the user is satisfied by the first relevant item. Across all values of $m$, $\mathrm{RR}$ dominates other metrics. $\mathrm{NDCG}$, often used for evaluating both top-$N$ recommendations and web search results, is the next more precision-oriented metric. AP, ‘recall at 1000’ ($\text{R}_{1000}$), and R-precision (RP) all have the same precision-orientation and are dominated by $\mathrm{NDCG}$. Rank-biased precision ($\mathrm{RBP}$) dominates $\mathrm{NDCG}$ and AP once we reach a modest number of relevant items. Both $\mathrm{RR}$ and $\mathrm{RBP}$ are not sensitive to the number of relevant items because neither is a function of $m$.
@@ 3.2. Total Search Efficiency
Although metrics such as AP are often referred to as ‘recall-oriented’, in this section, we focus on metrics that explicitly define recall as a construct. Such recall metrics for ranking systems come in two flavors. [^7] We exclude set-based metrics used in set-based retrieval or some technology-assisted review evaluation (Cormack and Grossman, 2016; Lewis et al., 2021), since they require systems to provide a cutoff in addition to a ranking. The first flavor of recall metrics measures the fraction of relevant items found after a user terminates their scan of the ranked list. Metrics like $\text{R}_{1000}$ and RP use a model of search depth to simulate how deep a user will scan. We can define exposure and normalization functions for $\text{R}_{k}$ and RP,
@@
Similarly, Zou and Kanoulas (2020) use ‘position of the last relevant item’ to evaluate high-recall tasks. By contrast, Rocchio (1964) proposed recall error, a metric based on the average rank of relevant items,
@@ 4. Robustness
In the context of a single ranking, we are interested in measuring its robustness in terms of its effectiveness for different possible users who might have issued the same request. [^8] Robustness in the information retrieval community has traditionally emphasized slightly different notions from our ranking-based perspective. For example, the TREC Robust track emphasized robustness of searcher effectiveness across information needs, focusing evaluation on difficult queries (Voorhees, 2004). Similarly, risk-based robust evaluation seeks to ensure that performance improvements across information needs are robust with respect to a baseline (Wang et al., 2012; Collins-Thompson et al., 2013). Meanwhile, Goren et al. (2018) proposed robustness as the stability of rankings across adversarial document manipulations. In the context of recommender systems, robustness has analogously focused on robustness of utility across users (Xu et al., 2020; Wen et al., 2022) and stability of rankings across adversarial content providers (Mobasher et al., 2007). In retrieval, this might occur when multiple users issue the same text-based query. In recommendation, this might occur when multiple users share the same engagement or session history, for example in cold-start or low-data situations. In both retrieval and recommendation, even a single user may have multiple intents depending on the context (e.g., an item may be relevant to a user when they are studying versus when they are relaxing). We present additional examples in Table 1 that demonstrate situations where data sparsity in the request can conceal two users with different intents. The concept of robustness across users or user-contexts is related to work in search engine auditing that empirically studies how effectiveness varies across different searchers issuing the same request (Mehrotra et al., 2017). In our work, we define robustness as the effectiveness of a ranking for the worst-off user (or user-context) who might have issued a request.
Table 1. Ambiguity in requests. In both retrieval and recommendation tasks, multiple users or the same user at different times may consider different items relevance.
request relevance
task implicit explicit user 1 user 2
web retrieval - query=‘salsa’ recipe dance
map retrieval location=‘NYC’ query=‘restaurant’ vegetarian Korean
movie recommendation watched=[‘Jaws’, ‘Piranha’] - nautical horror 1970s horror
music recommendation saved=[‘Black Sabbath’, ‘Bach’] query=‘study music’ metal classical
@@
Underlying our notion of robustness is a population-based perspective on ranking evaluation. Classic effectiveness measures can be interpreted as expected values over different user populations defined by different browsing behavior (Robertson, 2008; Sakai and Robertson, 2008; Carterette, 2011). For example, Robertson (2008) demonstrated that AP can be interpreted as the expected precision over a population of users with different recall requirements. More generally, Carterette (2011) demonstrated this for a large class of metrics. We can contrast this with online evaluation production environments where systems observe individual user behavior and do not need to resort to statistical models to capture different user behavior. So, just as recent work in the fairness literature disaggregates evaluation metrics to understand how performance varies across groups (Mehrotra et al., 2017; Ekstrand et al., 2018, 2022; Neophytou et al., 2022), we can disaggregate traditional evaluation metrics to understand how performance varies across implicit subpopulations of users or providers.
@@ 4.1. User Robustness
Given a ranking $\pi$, we would like to measure the worst-case effectiveness over a population of possible users.
@@ 4.1.1. Possible Information Needs
Similarly, in the context of recommendation, the notion of topical relevance can be interpreted as the potential utility—but not the psychological relevance or desirability—of an item to the user. So, while a particular item may be of interest to a user in general, at any one point in time, it may be undesired for any number of reasons. First, the context of a user can affect whether a particular item is desired. For example, in music recommendation, a user may be interested in bluegrass music in many contexts but, while studying, may prefer ambient music. Second, as with information retrieval, the desirability of a previously-consumed relevant item may degrade due to order effects. For example, in music recommendation, consider a user who listened to a specific relevant song $X$. There are two possible effects. On the one hand, satiation (Leqi et al., 2021) means that the user may not want to listen to $X$ immediately again, making it no longer relevant to that user. Or, obsessive listening (Ward et al., 2014; Conrad et al., 2019) means that the user may want to listen to $X$ over again, making other otherwise relevant songs no longer relevant to that user.
So, from the perspective of relevance, $\mathcal{R}$ should be considered the union of psychologically relevant items over all possible states of knowledge a user might have when approaching the system. Indeed, multiple authors describe topical relevance as necessary but not sufficient for psychological relevance (Boyce, 1982; Ruthven, 2014; Cooper, 1971). These papers suggest that rather than seeing a ranking system as acting to directly provide psychologically relevant items, it provides topically relevant items that are candidates to be scanned for psychologically relevant items by the user (Boyce, 1982). In light of this discussion, instead of considering any item in $\mathcal{R}$ as definitely satisfying the user’s information need, we consider it as only having a nonzero probability of satisfying the user’s information need.
While the concept of diversity (Castells et al., 2022; Santos et al., 2015), common to both recommendation and retrieval tasks, intends to support multiple information needs, they do not capture granular, item- or document-level needs. For example, measuring book recommendation diversity along the language dimension may detect poor performance for a group of users interested in Spanish language books, it will not be effective at detecting poor performance for Spanish language books unfamiliar to a particular user.
@@ 4.1.2. Robustness Across Possible Information Needs
In other words, the worst-off user is the one associated with $p_{m}$, the lowest-ranked relevant item. We present a proof in Appendix C.1. This result implies that recall-orientation captures the utility of a ranking for the worst-off user.
@@ 4.2.1. Possible Provider Preferences
Just as with information needs, each ranking consists of exposure of multiple possible providers. Consider the domains like job applicant ranking systems or dating platforms, where each item in the catalog is associated with an individual person. We assume that each relevant provider $i\in\mathcal{R}$ is interested in its cumulative positive exposure in the ranking (Section 2.3.2), $\eta(\pi,\mathcal{R},\{i\})$. In the more general case, providers can possibly be associated with multiple relevant items in $\mathcal{R}$. This might occur if a creator contributes multiple items to the catalog (e.g. multiple songs, videos, documents); or, a provider may aggregate content from multiple individual creators (e.g. publishers, labels). Even if we have metadata attributing groups of items to specific providers or creators, their preferences for exposure of those items may be unobserved. Provider preferences can themselves be complex, covering a broad set of commercial, artistic, and societal values (Hadida, 2015; Eikhof and Haunschild, 2006).
@@ 4.2.2. Robustness Across Possible Provider Preferences
Given this definition, we can show that $\operatorname{TSE}$ is equal to the utility of the worst-case provider (proof in Appendix C.1).
@@ 5. Lexicographic Evaluation
In Section 3, we defined recall-orientation from the perspective of users interested in finding the totality of relevant items and then proposed a new metric, $\operatorname{TSE}$, based on this interpretation. In Section 4, we demonstrated how $\operatorname{TSE}$ measures the worst-case utility for multiple definitions of users and providers, connecting it to notions of robustness and fairness, through Rawls’ difference principle. In this section, we will further develop the fairness perspective by combining recent work in preference-based evaluation with classic work in social choice theory, improving the nuance in worst-case analysis and allowing it to be useful as an evaluation tool. We will begin by discussing the practical limitations of $\operatorname{TSE}$ for evaluation (Section 5.1) before developing a preference-based evaluation method derived from social choice theory that generalizes $\operatorname{TSE}$ and improves its practical use (5.2).
@@ 5.1. Low Sensitivity of Total Search Efficiency
Although measuring worst-case performance and, as a result, Rawlsian fairness, $\operatorname{TSE}$ may not satisfy our desiderata for an evaluation method (Section 2.4). To understand why, consider two rankings $\pi$ and $\pi^{\prime}$ for the same request. When comparing a pair of systems, we are interested in defining a preference relation $\pi\succ\pi^{\prime}$. The worst-case preference $\pi\succ_{\text{WC}}\pi^{\prime}$, is defined as,
@@ 5.2. Lexicographic Recall
where $i=\max\{j\in[1\mathrel{{.}\,{.}}\nobreak m]:x_{i}\neq y_{j}\}$. We show an example of this process in Figure 9. This way of comparing rankings generates a total lexicographic ordering over vectors of the same dimensionality and is often used in the fairness literature to address ties when adopting Rawls’ difference principle. [^9] For further discussion of the connection between Rawls’ difference principle and leximin, see (Sen, 1976, 1977; Kolm, 2002; d’Aspremont and Gevers, 1977; Moulin, 2003).
@@
Although operating over $\mathcal{U}$ grounds our evaluation in possible user information needs, scoring and ranking $m^{+}$ subsets of $\mathcal{R}$ can be computationally intractable. So, just as we demonstrated that we only need to inspect the position of the last relevant item to compute $\succ_{\text{WC}}$, we can demonstrate that we only need to compare the rank positions of the relevant items to compute $\pi\succ_{\text{LR}}\pi^{\prime}$ (proof in Appendix C.2) and, therefore,
@@
where $i=\max\{j\in[1\mathrel{{.}\,{.}}\nobreak m]:p_{i}\neq p^{\prime}_{j}\}$. Moreover, although defined as a user-oriented metric, we can also demonstrate that this results in provider leximin as well (proof in Appendix C.2),
@@ 5.3. Sensitivity of Lexicographic Recall
We can demonstrate the higher sensitivity of lexirecall through simulation and analysis of the total space of permutations. In Figure 8, we demonstrated that, for a set of random paired rankings, the number of ties was high for traditional metrics and grew quickly for $\operatorname{TSE}$ as the cutoff $\tilde{n}$ decreased. Figure 8 also includes lexirecall, which exhibits substantially fewer ties than traditional metrics and $\operatorname{TSE}$. Independent of simulation, we are also interested in the probability of a tie over for randomly sampled pairs of complete rankings $\pi,\pi^{\prime}\in S_{n}$ (i.e. $\tilde{n}=n$). We can derive these probabilities (see Appendix E) as functions of $m$, $n$, and any parameters of the metric (e.g. $k$),
@@
To better understand the relationship between these probabilities, we display probabilities of ties for several retrieval depths in Table 3. Although Figure 8 demonstrated that $\operatorname{TSE}$ exhibited poor sensitivity when $\tilde{n}<n$, it is much more sensitive for complete rankings, in part because random complete rankings are less likely to share $p_{m}$ than imputed rankings. Both traditional recall metrics exhibit many more ties, especially as the retrieval depth grows. Amongst methods, lexirecall and AP demonstrate the few or no ties across corpus sizes. Table 4 presents the same results for varying numbers of relevant items. The results are consistent with Table 3, where traditional recall metrics exhibit a large number of ties, which decreases slowly with the number of relevant items. By comparison, $\operatorname{TSE}$, lexirecall, and AP show higher sensitivity with a negligible number of ties.